# Macro-from-Micro Planning for High-Quality and Parallelized Autoregressive Long Video Generation

### v2

@@ Abstract
Current autoregressive diffusion models excel at video generation but are generally limited to short temporal durations. Our theoretical analysis indicates that the autoregressive modeling typically suffers from temporal drift caused by error accumulation and hinders parallelization in long video synthesis. To address these limitations, we propose a novel planning-then-populating framework centered on Macro-from-Micro Planning (MMPL) for long video generation. MMPL sketches a global storyline for the entire video through two hierarchical stages: Micro Planning and Macro Planning. Specifically, Micro Planning predicts a sparse set of future keyframes within each short video segment, offering motion and appearance priors to guide high-quality video segment generation. Macro Planning extends the in-segment keyframes planning across the entire video through an autoregressive chain of micro plans, ensuring long-term consistency across video segments. Subsequently, MMPL-based Content Populating generates all intermediate frames in parallel across segments, enabling efficient parallelization of autoregressive generation. The parallelization is further optimized by Adaptive Workload Scheduling for balanced GPU execution and accelerated autoregressive video generation. Extensive experiments confirm that our method outperforms existing long video generation models in quality and stability. Generated videos and comparison results are in our project page.
Figure 1: We propose Macro-from-Micro Planning (MMPL), a novel paradigm for long-video generation. MMPL consistently delivers higher visual quality, faster generation speed, and stronger user preference than existing methods. Snapshots at 0s, 10s, 20s, and 30s (left) demonstrate the robustness of our method against temporal drift, which typically manifests as semantic shifts, color variations, and structural artifacts across long-horizon frames, while quantitative results showcase accelerated multi-GPU inference (top-right) and a dominant user preference rate (bottom-right).
@@ 1 Introduction
Long video generation is crucial for applications such as movie production (Polyak et al., 2024a; Zhao et al., 2025), virtual reality (Wu et al., 2025a; b), and digital human creation (Hu, 2024; Xiang et al., 2025; Zhang et al., 2025; Zhu et al., 2024). Despite significant advances in video synthesis, creating extended sequences with both temporal coherence and computational efficiency remains challenging (Ning et al., 2024).
@@
Analogous to the workflow of professional filmmakers, long video creation naturally benefits from a hierarchical plan-then-populate paradigm. In a typical movie production, the process does not proceed by shooting every frame in chronological order. Instead, the production team first develops a Macro Plan, a rough storyboard that captures the overall structure and key moments of the film. This Macro Plan consists of multiple Micro Plans, each representing an individual scene or shot. With this setup, different scenes can be filmed in parallel according to their Micro Plans, much like multiple crews shooting on separate sets at the same time. The Macro Plan then coordinates and assembles all these pieces into a coherent long movie. Such hierarchical planning improves the efficiency of film production while ensuring that the final movie remains seamless and coherent.
Building on this insight, we first perform a systematic analysis of error accumulation in AR and Non-AR video generation, revealing the fundamental mechanisms that drive long-term drift. Guided by these findings, we propose a novel plan-then-populate framework centered on Macro-from-Micro Planning (MMPL) for scalable, high-quality long video generation. MMPL operates via two complementary planning levels: Micro Planning efficiently predicts multiple keyframes of each segment simultaneously from its initial frame, capturing detailed local trajectories; Macro Planning autoregressively chains these segments by initializing each segment $S$ from the last keyframe of segment S1S-1, thus ensuring global narrative coherence across the entire video. Once all keyframes are established, MMPL-based Content Populating concurrently synthesizes intermediate frames between keyframes within each segment, adhering to boundary constraints and eliminating sequential frame dependencies. To further optimize pipeline efficiency, we introduce an adaptive workload scheduling strategy that dynamically allocates GPU resources. This approach significantly reduces the overall generation time to approximately one-third of the original, without relying on distillation-based acceleration, while preserving high visual fidelity.
@@ 2 Related Work
Planning Prediction. A closely related work, FramePack-Plan (Zhang & Agrawala, 2025), mitigates error accumulation via step-wise frame jumping, and compresses context to extend video length. In contrast, our Macro-from-Micro framework introduces three key innovations. First, we adopt a two-level hierarchical planning scheme: a Micro Plan predicts segment-level keyframes, and a Macro Plan, composed of overlapping Micro Plans, forms a coherent global storyline through autoregressive scheduling. Second, each Micro Plan produces all pre-planned keyframes for its segment in a single forward pass conditioned only on the initial frame, drastically compressing the autoregressive chain. Finally, once the full Macro Plan is obtained, the remaining content within all segments is synthesized in parallel, achieving high throughput while preserving local temporal coherence.
3 Drift Analysis in Long Video Generation
$p_{\theta}(x)=\prod_{t=1}^{T}p_{\theta}(x^{t}\mid x^{<t}),$ (1)
$\mathcal{R}(\pi_{BC})=L^{I}(\pi_{BC})-L^{I}(o).$ (3)
$T\epsilon\leq\mathcal{R}(\pi_{BC})\leq T^{2}\epsilon,$ (4)
Non-Autoregressive (Non-AR) Models. Non-AR models generate future frames jointly in a single forward pass, conditioned solely on the initial frame. Formally, the joint distribution over frames $x^{2:i}=\{x^{2},\dots,x^{i}\}$ is modeled as $p_{\theta}(x^{2:i}\mid x^{1})$, allowing all frames to be optimized together rather than sequentially. Such joint prediction eliminates step-wise dependency, thereby preventing the cumulative error propagation inherent in autoregressive models and offering greater stability for long-horizon synthesis, as exemplified in image-to-video generation.
@@ 4.1 Macro-from-Micro Planning
Motivated by the analysis in Sec. 3, we observe that autoregressive models accumulate errors proportionally to the number of propagation steps, whereas non-autoregressive models decouple errors from the step count through joint optimization. To exploit the complementary strengths of both paradigms, we introduce Macro-from-Micro Planning (MMPL), a unified planning method comprising two key components: Micro-Planning and Macro-Planning.
@@ 5 Experiments
Training Details. We implement MMPL on Wan2.1-T2V-14B (Wang et al., 2025a), a bidirectional DiT-based (Peebles & Xie, 2023) Flow Matching model originally designed for 5-second video generation. To enable efficient long-horizon modeling, FlexAttention (Dong et al., 2024) is adopted during training for scalable attention, and FlashAttention-v3 (Dao et al., 2022) is employed during inference to further accelerate sampling. The model is fine-tuned on 50,000 manually curated high-quality videos at $832{\times}480$ resolution, providing diverse and clean training data that supports stable optimization and long-horizon generation. Training is conducted for 8,000 iterations on 32 H100 GPUs using the AdamW optimizer with a learning rate of $1\times 10^{-5}$. For hierarchical planning, we set $t_{a}=2,3,4$, $t_{b}=9,10,11$, and $t_{c}=19,20,21$, corresponding to early, midpoint, and late planning frames that guide the segment-wise generation process. Additional hyperparameters and ablation settings are provided in the supplementary material.
Evaluation. We evaluate on VBench-long benchmark (Huang et al., 2024; Zheng et al., 2025), which measures subject consistency, background consistency, motion smoothness, aesthetic quality, and imaging quality, jointly capturing temporal stability and perceptual fidelity. For the primary study, we generate 30s videos for 120 randomly sampled MovieGen (Polyak et al., 2024b) prompts on a single H100 GPU. We further perform a user study to complement the quantitative metrics. Specifically, for each baseline, 19 videos of approximately 30 s are generated using the first 19 MovieGen prompts, and 29 independent participants are recruited to perform pairwise comparisons, selecting the video that better reflects the input prompt in terms of visual quality and semantic fidelity. This combination of objective and subjective assessments provides a rigorous evaluation of both numerical performance and perceptual quality. Further details of the user study are provided in the supplementary material.
Table 1: Evaluation metrics for the other baselines and our planning-based diffusion models. The first five metrics are automatically computed by VBench, while the last three are obtained through human evaluation, providing complementary assessments of perceptual quality and long-horizon coherence beyond the automatic metrics.
@@
Figure 8: Qualitative comparisons of different MMPL variants.
Ablation Studies. In long-range video generation, the placement of Planning frames within each segment during Micro Planning plays a pivotal role in the overall performance of MMPL, as it critically impacts both temporal smoothness and structural consistency. To validate this hypothesis, we conduct an ablation study with three Micro Planning variants: (i) Planning without early frames, omitting frames near the initial timestep; (ii) Planning without the midpoint frame, removing the central anchor; and (iii) the full MMPL strategy, retaining all Planning frames. As shown in Table 6, our complete MMPL configuration consistently outperforms the other two variants across all evaluation metrics. Furthermore, as shown in Figure 8, we conduct qualitative comparisons among these variants. The results clearly demonstrate that the full MMPL strategy not only produces smoother temporal transitions and more stable long-horizon content, but also, in contrast, the other variants—due to the absence of certain planning frames—tend to exhibit discontinuous transitions or noticeable jumps in the predicted frames at the corresponding positions.
6 Discussion
Table 2: Compact ablation study on Micro Planning setups and training strategies.
@@ 6 Discussion
[…]
Compatibility with Acceleration and Distillation Methods. Our paradigm is inherently compatible with acceleration techniques such as DMD and other distillation methods, requiring no modifications to the original architecture. During training, we simply adjust the attention mask to control the visible frame range, while at inference we achieve efficient generation by reorganizing the decoding order of video segments. This inherent compatibility allows Macro-from-Micro to seamlessly integrate with existing acceleration pipelines. Incorporating model distillation or similar strategies in future work may further enhance inference efficiency while preserving generation quality.
Compatibility with Self Forcing Approaches. Macro-from-Micro is also complementary to self-correcting strategies that mitigate step-wise autoregressive errors, such as Self Forcing. During training, the model typically predicts the next frame by denoising conditioned on ground-truth video frames. Replacing these ground-truth frames with previously generated predictions naturally transitions the model into a Self-Macro-from-Micro regime. This hybrid strategy can further extend the duration of generable videos and substantially improve temporal consistency over long sequences.
Limitations. Although Macro-from-Micro Planning substantially mitigates the accumulation of prediction errors, slight quality degradation may still arise in ultra-long video generation. Because the content frames within each segment are interpolated between planning frames, motion continuity can weaken near boundaries. In addition, the substantial temporal span of long videos means that a single text prompt often aligns only with the early content and fails to capture the complete video semantics. Since our current framework does not dynamically update prompts during generation, this static conditioning leads to repetitive or even collapsed content in later segments.

### v3

@@ Abstract
Current autoregressive diffusion models excel at video generation but are generally limited to short temporal durations. Our theoretical analysis indicates that the autoregressive modeling typically suffers from temporal drift caused by error accumulation and hinders parallelization in long video synthesis. To address these limitations, we propose a novel planning-then-populating framework centered on Macro-from-Micro Planning (MMPL) for long video generation. MMPL sketches a global storyline for the entire video through two hierarchical stages: Micro Planning and Macro Planning. Specifically, Micro Planning predicts a sparse set of future keyframes within each short video segment, offering motion and appearance priors to guide high-quality video segment generation. Macro Planning extends the in-segment keyframes planning across the entire video through an autoregressive chain of micro plans, ensuring long-term consistency across video segments. Subsequently, MMPL-based Content Populating generates all intermediate frames in parallel across segments, enabling efficient parallelization of autoregressive generation. The parallelization is further optimized by Adaptive Workload Scheduling for balanced GPU execution and accelerated autoregressive video generation. Extensive experiments confirm that our method outperforms existing long video generation models in quality and stability. Generated videos and comparison results are in the Demo page.
Figure 1: We propose Macro-from-Micro Planning (MMPL), a paradigm for long-video generation that achieves higher visual quality, faster speed, and stronger user preference than existing methods. Snapshots at 0s, 10s, 20s, and 30s (left) show robustness against temporal drift—semantic shifts, color changes, and structural artifacts—while quantitative results highlight accelerated multi-GPU inference (top-right) and dominant user preference (bottom-right).
@@ 1 Introduction
Long video generation is crucial for applications such as movie production (Polyak et al., 2024; Zhao et al., 2025), virtual reality (Wu et al., 2025a; b), and digital human creation (Hu, 2024; Xiang et al., 2025; Zhang et al., 2025; Zhu et al., 2024). Despite significant advances in video synthesis, creating extended sequences with both temporal coherence and computational efficiency remains challenging (Ning et al., 2024).
@@
Analogous to the workflow of professional filmmakers, long video creation naturally benefits from a hierarchical plan-then-populate paradigm. In a typical movie production, the process does not proceed by shooting every frame in chronological order. Instead, the production team first develops a Macro Plan, a rough storyboard that captures the overall structure and key moments of the film. This Macro Plan consists of multiple Micro Plans, each representing an individual scene or shot. With this setup, different scenes can be filmed in parallel according to their Micro Plans, much like multiple crews shooting on separate sets at the same time. The Macro Plan then coordinates and assembles all these pieces into a coherent long movie. Such hierarchical planning improves the efficiency of film production while ensuring that the final movie remains seamless and coherent. Building on this insight, we first perform a systematic analysis of error accumulation in AR and Non-AR video generation, revealing the fundamental mechanisms that drive long-term drift. Guided by these findings, we propose a novel plan-then-populate framework centered on Macro-from-Micro Planning (MMPL) for scalable, high-quality long video generation. MMPL operates via two complementary planning levels: Micro Planning efficiently predicts multiple keyframes of each segment simultaneously from its initial frame, capturing detailed local trajectories; Macro Planning autoregressively chains these segments by initializing each segment $S$ from the last keyframe of segment S1S-1, thus ensuring global narrative coherence across the entire video. Once all keyframes are established, MMPL-based Content Populating concurrently synthesizes intermediate frames between keyframes within each segment, adhering to boundary constraints and eliminating sequential frame dependencies. To further optimize pipeline efficiency, we introduce an adaptive workload scheduling strategy that dynamically allocates GPU resources. This approach significantly reduces the overall generation time to approximately one-third of the original, without relying on distillation-based acceleration, while preserving high visual fidelity.
@@ 2 Related Work
Planning Prediction. A closely related work, FramePack-Plan (Zhang & Agrawala, 2025), mitigates error accumulation via step-wise frame jumping, and compresses context to extend video length. In contrast, our Macro-from-Micro framework introduces three key innovations. First, we adopt a two-level hierarchical planning scheme: a Micro Plan predicts segment-level keyframes, and a Macro Plan, composed of overlapping Micro Plans, forms a coherent global storyline through autoregressive scheduling. Second, each Micro Plan produces all pre-planned keyframes for its segment in a single forward pass conditioned only on the initial frame, drastically compressing the autoregressive chain. Finally, once the Macro Plan is obtained, the remaining content within all segments is synthesized in parallel, achieving high throughput while preserving temporal coherence.
@@ 3.1 Macro-from-Micro Planning
Motivated by the analysis in the supplementary material, we observe that autoregressive models accumulate errors proportionally to the number of propagation steps, whereas non-autoregressive models decouple errors from the step count through joint optimization. To exploit the complementary strengths of both paradigms, we introduce Macro-from-Micro Planning (MMPL), a unified planning method comprising two key components: Micro-Planning and Macro-Planning.
@@ 4 Experiments
Training Details. We implement MMPL on Wan2.1-T2V-14B (Wang et al., 2025a), a DiT-based (Peebles & Xie, 2023) Flow Matching model originally built for 5-second videos. To support long-horizon modeling, we adopt FlexAttention (Dong et al., 2024) for scalable training and FlashAttention-v3 (Dao et al., 2022) for fast inference. The model is fine-tuned on 50k curated high-quality videos at $832{\times}480$ resolution, ensuring diverse and clean data for stable optimization. Training runs for 8,000 iterations on 32 H100 GPUs with AdamW at a $1\times 10^{-5}$ learning rate. For hierarchical planning, we set $t_{a}=2,3,4$, $t_{b}=9,10,11$, and $t_{c}=19,20,21$, corresponding to early, midpoint, and late planning frames guiding segment-wise generation. Additional hyperparameters and ablation settings are provided in the supplementary material.
Evaluation. We evaluate on the VBench-long benchmark (Zheng et al., 2025), which assesses subject and background consistency, motion smoothness, aesthetic quality, and imaging quality, jointly reflecting temporal stability and perceptual fidelity. For the main study, we generate 30s videos from 120 randomly sampled MovieGen (Polyak et al., 2024) prompts on a single H100 GPU. To complement these quantitative metrics, we also conduct a user study: for each baseline, 19 videos of about 30 s are generated using the first 19 MovieGen prompts, and 29 participants perform pairwise comparisons, selecting the video that better matches the prompt in terms of visual quality and semantic fidelity. This combination of objective metrics and human judgments provides a rigorous evaluation of both numerical performance and perceptual quality. Details of the user study are provided in the supplementary material.
Table 1: Evaluation metrics for the other baselines and MMPL. The first five metrics are automatically computed by VBench, while the last three are obtained through human evaluation.
@@
Figure 8: Qualitative comparisons of different MMPL variants.
@@
Table 2: Ablation studies on planning setups.
@@
Table 3: Ablation studies on training strategies.
VBench
Variant Subj. Back. Mot. Aes. Img.
Training Strategy
Freeze 0.838 0.923 0.973 0.484 0.503
Only Q,K 0.970 0.962 0.987 0.612 0.647
Only Self-Attention 0.980 0.968 0.992 0.628 0.661
Ablations on Micro-Planning Frame Placement. The placement of Planning frames within each segment during Micro Planning is pivotal for MMPL, shaping temporal and structural consistency. We validate this via an ablation with three variants: (i) without early frames (omit frames near the start); (ii) without the midpoint frame (remove the central anchor); and (iii) the full MMPL that retains all Planning frames. As shown in Table 3, the full configuration leads across all metrics. Qualitatively (Figure 8), it yields smoother transitions and more stable long-horizon content, whereas the ablated variants exhibit discontinuities or noticeable jumps around the missing frames.
Ablations on Model Training Strategy. We compare three update policies for the video generation model: Freeze freezes all parameters; Only Q,K updates only the self-attention query and key projections; Only Self-Attention updates Q, K, V and the attention output, while feed-forward layers and embeddings remain frozen. As shown in Table 3, updating the whole self-attention yields the best scores across all metrics. Training only Q,K is lighter but slightly weaker. Freezing performs worst and shows larger temporal drift and inconsistency.
5 Discussion
Compatibility with Acceleration and Distillation Methods. Our paradigm is naturally compatible with acceleration techniques such as DMD and other distillation approaches, requiring no architectural changes. During training, we only adjust the attention mask to control the visible frame range, while at inference efficiency is improved by reorganizing the decoding order of video segments. This flexibility allows Macro-from-Micro to plug into existing acceleration pipelines. The results of adapting model distillation and related strategies are provided in the supplementary material.
Compatibility with Self-Forcing Approaches. MMPL also complements self-correcting strategies that mitigate step-wise autoregressive errors, such as Self Forcing. In standard training, the model predicts the next frame by denoising conditioned on ground-truth video frames; replacing these with previously generated predictions naturally yields a Self-MMPL regime. This hybrid setup extends the duration of generable videos and improves temporal consistency over long sequences.
Limitations. Although Macro-from-Micro Planning substantially mitigates the accumulation of prediction errors, generating hour-level videos remains challenging due to the reliance on a single static text prompt. The substantial temporal span of long videos means that a single text prompt often aligns only with the early content and fails to capture the full video semantics. Since our current framework does not dynamically update prompts during generation, this static conditioning limits the ability to produce coherent and continuous content across longer segments.
@@ 6 Conclusion
Appendix A Data Preparation
Data Sources and Filtering. Our dataset comprises two components: (i) licensed commercial videos purchased from authorized providers; and (ii) web videos manually collected from open platforms, primarily Mixkit, Pexels, and Pixabay. For all candidate videos, we first generate textual captions with Qwen-72B and compute an aesthetic score using the LAION aesthetic predictor. For licensed data, we rank by the aesthetic score and retain only the top 1%. For web-collected data, we conduct human quality control to remove low-quality clips, content–caption mismatches, and potential copyright risks, and to verify caption consistency. We then merge the two subsets, yielding approximately $\sim$50k high-quality samples to train our long-video generation model.
Structured Annotation Pipeline. To obtain rich and structured annotations, we drive Qwen-72B with carefully designed instruction prompts to analyze video frames and output a JSON object with a fixed schema. The JSON includes: a short scene summary (short_caption); a dense contextual description (dense_caption, covering main subject, background, visual style, camera movement, shot type, lighting, and atmosphere); detailed subject descriptions (for persons: facial expressions, emotional state, and ethnicity); background information; standardized style/shot/movement labels; aesthetic tags; and role statistics (e.g., number of humanoid characters, coverage extent, depiction style, and motion dynamics). Concretely, short_caption is generated with the instruction ”Brief scene summary in 1 sentence”, while dense_caption uses ”Detailed context including main subject, background, visual style, camera movement tech, shot type, lighting, and atmosphere”. All outputs are in English, follow predefined field orders and constraints, and employ standardized vocabulary for key attributes.
Quality Control and Training Setup. After annotation, we re-evaluate aesthetic quality with the LAION predictor to ensure consistency. During training, for each of the $\sim$50k videos we sample the conditioning text with probability 0.8 from dense_caption and with probability 0.2 from short_caption. This strategy preserves the high information density of dense captions while maintaining robustness and diversity from concise summaries. As shown in Figure 9, we present several representative examples from our curated dataset.
Figure 9: Examples of training samples. The dataset combines licensed and web-collected videos, curated via aesthetic scoring and manual screening.
Appendix B User Study Details
To complement the quantitative metrics, we conduct a user study on long-video generation. In each trial, participants evaluate five videos generated from the same prompt by ranking them (1 = best, 5 = worst) along three dimensions: text–visual alignment, content consistency, and long-sequence color stability. In addition, participants select a single overall favorite video.
Figure 10: User study instruction screenshots.
This protocol provides fine-grained human judgments on both quality and temporal robustness that are not fully captured by automated metrics. Detailed instructions are shown in Figure 10.
Appendix C Training Settings
C.1 Hyperparameter Settings
Most experiments are conducted on 32 NVIDIA GPUs (80 GB each), using a per-GPU batch size of 1 without gradient accumulation. The detailed hyperparameters are summarized in Table 4. Training the Teacher Forcing 14B model for 8,000 steps required about three days, while the DMD 1.3B model reached 8,000 steps within roughly one day.
Table 4: Specification of training hyperparameters
Hyperparameters Teacher Forcing Self Forcing
Generate network Wan2.1-T2V-14B Wan2.1-T2V-1.3B
Real score network N/A Wan2.1-T2V-14B
Fake score network N/A Wan2.1-T2V-14B
Batch size 32 32
Optimizer ($G_{\theta}$) AdamW, $\beta_{1}=0$, $\beta_{2}=0.999$, $\epsilon=1\!\times\!10^{-8}$, weight_decay =0.01=0.01 Adam, $\beta_{1}=0$, $\beta_{2}=0.999$, $\epsilon=1\!\times\!10^{-8}$, weight_decay =0.01=0.01
Optimizer ($f_{\psi}$) N/A Adam, $\beta_{1}=0$, $\beta_{2}=0.999$, $\epsilon=1\!\times\!10^{-8}$, weight_decay =0.01=0.01
Learning rate ($G_{\theta}$) $1\!\times\!10^{-5}$ $2\!\times\!10^{-6}$
Learning rate ($f_{\psi}$) N/A $4\!\times\!10^{-7}$
Gen./Cri. update ratio N/A 5
EMA decay N/A 0.99
C.2 Planning Settings
Settings. To clarify the generation process, we detail the model’s computation using the 21 latent tokens of the full Wan model as shown in Figure 11. Tokens are indexed from 0: indices 0–1 correspond to the initial frame, indices 2–3 to early planning frames, indices 10–12 to midpoint planning frames, and indices 19–20 to terminal planning frames.
Figure 11: Overview of our planning-based inference on an 81-frame sequence with 21 latent tokens (full Wan model). Pre-Planning latent tokens at the beginning, midpoint, and terminal positions serve as stable anchors in the denoising schedule, guiding the synthesis of all intermediate frames and ensuring long-range temporal coherence.
Analysis. While the proposed planning setup places anchors over long horizons, a central challenge remains: enabling the autoregressive (AR) decoder to effectively exploit these anchors when synthesizing intermediate frames. Relying on a single planning token at the early, midpoint, and terminal boundaries is intrinsically fragile in the presence of the AR decoder’s pronounced recency bias—the tendency to overweight the most recent observations while underutilizing distant context. This bias causes the model, at each sub-segment junction, to condition predominantly on the tail of the preceding sub-segment, thereby inheriting and amplifying residual errors and inducing cross-boundary propagation. Consequently, a single planning token per boundary is insufficient to arrest drift arising from accumulated errors. Formally, this bias in MMPL is expressed in Eq. 6:
$p\big(x^{s_{k}+1:e_{k}-1}\mid{x}^{1:s_{k}},x^{e_{k}}\big)\approx p\big(x^{s_{k}+1:e_{k}-1}\mid x^{s_{k}-K:s_{k}},x^{e_{k}}\big).$ (6)
Here, $s_{k}$ and $e_{k}$ denote the starting and ending reference indices of sub-segment $k$, corresponding to the pre-planned planning frames. The hyperparameter $K$ specifies the size of the recent-context window on which the AR decoder conditions—namely, the $K$ frames immediately preceding $s_{k}$. Because $\{s_{k}-K:s_{k}\}$ overlaps with the tail of the previous sub-segment, residual errors inevitably leak into the current one, leading to error propagation across boundaries.
To counteract this bias, we replace the single predecessor at each boundary with a local multi-frame set. Concretely,
$\mathcal{P}_{s_{1}}=\{2,3\},\quad\mathcal{P}_{e_{1}}=\{10,11,12\},\quad\mathcal{P}_{s_{2}}\approx\{10,11,12\},\quad\mathcal{P}_{e_{2}}=\{19,20\},$
where $\mathcal{P}_{s_{k}}$ denotes the local index set of expanded pre-planning frames around boundary $s_{k}$. Using these expanded anchors, the conditional distribution for sub-segment $k$ is refined to
$p\big(x^{s_{k}+1:e_{k}-1}\mid{x}^{1:s_{k}},x^{e_{k}}\big)\approx p\big(x^{s_{k}+1:e_{k}-1}\mid x^{(s_{k}-2-K):(s_{k}-3)},\,x^{s_{k}-2:s_{k}},\,x^{e_{k}}\big).$ (7)
Conditioning on a compact bundle of early-step, low-drift frames—rather than a single predecessor—dilutes residual errors inherited from the previous sub-segment. At the same time, the model’s recency bias naturally prioritizes the most recent elements within this bundle, thereby stabilizing long-horizon synthesis and suppressing cross-boundary error propagation without discarding information from the planned anchors.
Appendix D Error Accumulation Analysis in AR models
$p_{\theta}(x)=\prod_{t=1}^{T}p_{\theta}(x^{t}\mid x^{<t}),$ (8)
$\mathcal{R}(\pi_{BC})=L^{I}(\pi_{BC})-L^{I}(o).$ (10)
$T\epsilon\leq\mathcal{R}(\pi_{BC})\leq T^{2}\epsilon,$ (11)
Appendix E Importance of VAE
We compare the extrapolation procedure from the public Causvid (Yin et al., 2025) codebase against our proposed drift-resilient re-encoding and decoding strategy as shown in Figure 12. When extrapolation goes beyond the training context length and requires segment stitching, the baseline suffers from severe color drift and visual artifacts, whereas our method effectively mitigates these degradations.
Figure 12: Qualitative comparisons on video extrapolation.
Appendix F Noise Initialization Strategy for Smoothing Generation
In this work, we propose a specialized noise initialization strategy to address potential temporal discontinuities and instability at the transition boundaries between planning frames and content frames as shown in Figure 13. This approach ensures smooth visual transitions by strategically incorporating noise information from adjacent planning frames during the content frame generation process. Let $P_{n-1}$ and $P_{n}$ represent the planning frames at temporal positions n1n-1 and $n$, respectively, and let $C_{n+1}$ denote the target content frame at position $n+1$. To establish the theoretical foundation, we first recall the standard diffusion forward process formulation. Given a clean frame $\mathbf{x}_{0}$ at diffusion timestep $t$, the noisy observation $\mathbf{x}_{t}$ is generated through the Gaussian perturbation:
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}{t}}\mathbf{x}_{0},(1-\bar{\alpha}{t})\mathbf{I}),$ (13)
where $\bar{\alpha}_{t}$ denotes the cumulative product of the noise schedule coefficients. This process can be equivalently expressed as:
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}{t}}\cdot\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}{t}}\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (14)
Building upon this formulation, our methodology initializes the noise vector $\epsilon_{C_{n+1}}$ for the content frame $C_{n+1}$ through a weighted interpolation of the noise vectors associated with the preceding planning frames. Specifically, the initialization follows:
$\epsilon_{C_{n+1}}=\alpha\cdot\epsilon_{P_{n}}+(1-\alpha)\cdot\epsilon_{P_{n-1}},$ (15)
where $\epsilon_{C_{n+1}}$ represents the noise vector utilized in the reverse diffusion process for generating content frame $C_{n+1}$, $\epsilon_{P_{n}}$ and $\epsilon_{P_{n-1}}$ correspond to the noise vectors derived from planning frames $P_{n}$ and $P_{n-1}$ This noise initialization strategy ensures a continuous evolution of stochastic patterns across frame boundaries, effectively mitigating visual artifacts and temporal inconsistencies. By controlling the interpolation weight $\alpha$, our method provides precise adjustment over the temporal smoothness characteristics, enabling stable and coherent video generation while maintaining high visual quality throughout the sequence.
Figure 13: Framework for stable and smoothing frame generation via coherent noise initialization.
Appendix G Scalability
G.1 Image-to-Video Extension
Our framework is not restricted to the text-to-video (T2V) task; it can be seamlessly extended to image-to-video (I2V) generation without introducing any architectural modifications or additional image encoders. This flexibility derives from the unified autoregressive design, which only requires lightweight adjustments to the number and ordering of autoregressive steps. As a result, the framework adapts naturally to different input modalities while maintaining temporal consistency and generation quality as shown in Figure 14.
Figure 14: Qualitative results of extending our unified autoregressive framework from text-to-video (T2V) to image-to-video (I2V) generation. Without any architectural modifications or additional image encoders, the framework adapts seamlessly by only adjusting the number and ordering of autoregressive steps, while preserving temporal consistency and visual quality.
G.2 Adaptation to Self-Forcing and DMD
Our approach can be seamlessly integrated with self-forcing strategies without any architectural modifications. Specifically, it only requires adjusting the attention visibility range and the prediction order during both training and inference. This lightweight adaptation enables direct compatibility with existing self-forcing pipelines, while retaining the benefits of our planning-based design. Combined with parallelized decoding, the resulting system achieves substantial inference speedups, sustaining over 32 FPS in long-horizon video generation as shown in Figure 15.
Figure 15: Integration of our framework with Self Forcing (Huang et al., 2025a) and DMD (Yin et al., 2025) strategies. The adaptation requires no architectural changes—only modifications to the attention visibility range and prediction order during training and inference. Combined with parallelized decoding, the method achieves substantial inference acceleration, sustaining over 32 FPS in long-horizon video generation.
Appendix H More Qualitative Results
To better demonstrate the robustness of our model, we present additional experimental results on 30s long video generation, as shown in Figure 16.
Figure 16: Additional qualitative results of 30s long video generation. Our model produces temporally coherent and visually consistent sequences across diverse scenarios, further demonstrating its robustness.